OpenStudy (anonymous):

A number is decreased by seven. Then, the new number is multiplied by 3 to get an answer of -9. What is the original number? -30 30 4 -4

OpenStudy (anonymous):

It's 4

OpenStudy (anonymous):

4-7 = -3 *3 = -9

OpenStudy (mathstudent55):

A number: x decreased by 7: x - 7 multiplied by 3: 3(x - 7) answer is -9: 3(x - 7) = -9 Solve for x.

OpenStudy (anonymous):

Let the number=x 3(x-7)=-9 solve for x